# Persistent Homology for High-dimensional Data Based on Spectral Methods

### v2

@@ Appendix I Hyperparameter sensitivity
Figure S28: Void detection score for persistent homology with various distances on a noisy torus in ambient $\mathbb{R}^{50}$. Methods relying on 2D embeddings did not find the void for any noise level. Only $t$-SNE graph and UMAP graph could reliably improve above the Euclidean distance and only for low noise levels. However, they had unstable behavior for higher noise levels, resulting in high uncertainties. We suspect that a higher sampling density would benefit effective resistance and diffusion distance a lot, but the computational complexity of persistent homology makes such experiments difficult.
Figure S29: Detection scores for all hyperparameter settings for all six single-cell datasets. We omitted DTM as no setting passed the thresholding on any dataset. The black bar refers to correlation distance on the Malaria dataset and to Euclidean distance on the others. Extension of Figure 9. $t$-SNE graph and UMAP graph can perform very well, but are very hyperparameter-dependent. Their embedding variants often perform well, but collapse on some datasets. The spectral behave similarly, but perform better on average.
Figure S30: Detection scores for effective resistance, Euclidean distance, and various $\ell_{p}$ distances on the noisy circle, interlinked circles, and the eyeglasses dataset in ambient $\mathbb{R}^{50}$. For $p>2$ the $\ell_{p}$ distance becomes non-isotropic which outweighed any positive effect of accumulating less of the noise. The detection scores were on par with or worse than those of the Euclidean distance.

### v3

@@ Appendix I Details on the distances used in our benchmark
Figure S32: Diffusion distances failed on the torus with $n=1000$ because its eigenvalue decay suppressed the relevant eigenvectors. a. 2-loop detection score for the torus in $d=50$ ambient dimensions. Diffusion distances with $t=2$ diffusion steps were on par with effective resistance and Euclidean distance. b. Decay of eigenvalues in various spectral distances on the noiseless torus. Diffusion distances with $t=8,64$ only had contribution below $0.1$ for the fifth and sixth eigenvectors, while effective resistance and diffusion distance with $t=2$ had substantial contributions from the first ${\sim}10$ eigenvectors. c – f. Eigenvectors of the symmetric graph Laplacian of a symmetric $100$-nearest-neighbor graph of the noiseless torus. Coordinates are the angles of each point along ($\phi$) and around ($\theta$) the tube of the torus. The loop along the tube is encoded in the first two eigenvectors, the loop around the tube in the fifth and sixth eigenvectors.
Figure S33: Void detection score for persistent homology with various distances on a noisy torus in ambient $\mathbb{R}^{50}$. Methods relying on 2D embeddings did not find the void for any noise level. Only $t$-SNE graph and UMAP graph could reliably improve above the Euclidean distance and only for low noise levels. However, they had unstable behavior for higher noise levels, resulting in high uncertainties. We suspect that a higher sampling density would benefit effective resistance and diffusion distance (as we saw for loop detection in Figure S31), but the computational complexity of persistent homology makes such experiments difficult.
Figure S34: Detection scores for all hyperparameter settings for all six single-cell datasets. We omitted DTM as no setting passed the thresholding on any dataset. The black bar refers to correlation distance on the Malaria dataset and to Euclidean distance on the others. Extension of Figure 10. $t$-SNE graph and UMAP graph could perform very well, but were very hyperparameter-dependent. Their embedding variants often performed well, but collapsed on some datasets. The spectral methods behaved similarly, but on average performed better.